# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v1

@@ Abstract
Training deep neural networks—and more recently, large models—demands efficient and scalable optimizers. Adaptive gradient algorithms like Adam, AdamW, and their variants have been central to this task. Despite the development of numerous variance reduction algorithms in the past decade aimed at accelerating stochastic optimization in both convex and nonconvex settings, variance reduction has not found widespread success in training deep neural networks or large language models. Consequently, it has remained a less favored approach in modern AI. In this paper, to unleash the power of variance reduction for efficient training of large models, we propose a unified optimization framework, MARS (Make vAriance Reduction Shine), which reconciles preconditioned gradient methods with variance reduction via a scaled stochastic recursive momentum technique. Within our framework, we introduce three instances of MARS that leverage preconditioned gradient updates based on AdamW, Lion, and Shampoo, respectively. We also draw a connection between our algorithms and existing optimizers. Experimental results on training GPT-2 models indicate that MARS consistently outperforms AdamW by a large margin.
@@ 1 Introduction
Adaptive gradient methods such as Adam (Kingma, 2014) and AdamW (Loshchilov, 2017) have become the predominant optimization algorithms in deep learning. With the surge of large language models, the majority of the renowned models, including GPT-2 (Radford et al., 2019), GPT-3 (Brown, 2020), PaLM (Chowdhery et al., 2023) and Llama 3 (Dubey et al., 2024) are trained with adaptive gradient methods. Numerous efforts have been made to improve adaptive gradient methods from both first-order and second-order optimization perspectives. For example, You et al. (2019) introduced LAMB, a layerwise adaptation technique that boosts training efficiency for BERT (Devlin, 2018). Using symbolic search, Chen et al. (2023) developed Lion, achieving faster training and reduced memory usage. Liu et al. (2023) designed Sophia, leveraging stochastic diagonal Hessian estimators to accelerate training. Gupta et al. (2018) proposed Shampoo, which performs stochastic optimization over tensor spaces with preconditioning matrices for each dimension. Anil et al. (2020) further refined Shampoo to give a scalable, practical version. Recently, Vyas et al. (2024) showed that Shampoo is equivalent to Adafactor (Shazeer and Stern, 2018) in the eigenbasis of Shampoo’s preconditioner, and introduced SOAP, which stabilizes Shampoo with Adam. Importantly, recent studies (Kaddour et al., 2024; Zhao et al., 2024) have shown that these optimizers perform on par with AdamW in LLM pretraining, yet do not outperform it. This suggests the ongoing challenge in developing adaptive gradient methods superior to Adam and AdamW for large-scale model training.
Since adaptive gradient methods face challenges of high stochastic gradient variance, and language model training inherently involves a high-variance optimization problem (McCandlish et al., 2018), it is natural to consider variance reduction techniques to address this challenge. There exists a large body of literature on variance reduction for stochastic optimization, such as SAG (Roux et al., 2012), SVRG (Johnson and Zhang, 2013) and STORM (Cutkosky and Orabona, 2019), which can improve the convergence of stochastic optimization. However, variance reduction has not found widespread success in training deep neural networks or large language models. Defazio and Bottou (2019) discussed why variance reduction can be ineffective in deep learning due to factors such as data augmentation, batch normalization, and dropout, which disrupt the finite-sum structure required by variance reduction principles. Nevertheless, in training language models, data augmentation, batch normalization, and dropout are nowadays rarely used, which opens the door for applying variance reduction techniques in optimizing these models. This naturally leads to the following research question:
@@
In this paper, we answer the above question affirmatively by introducing a novel optimization framework called MARS (Make vAriance Reduction Shine), which incorporates variance reduction into adaptive gradient methods. Notably, we introduce a scaling parameter into the stochastic recursive momentum (STORM) (Cutkosky and Orabona, 2019) to adjust the strength of variance reduction and define a new gradient estimator. This gradient estimator undergoes gradient clipping and is subsequently subjected to exponential averaging. When the variance reduction strength is set to $1$, it recovers the vanilla STORM momentum. In addition, the second-order momentum update is defined by the reweighted intermediate variable. These together ensure optimization stability throughout the training process. We summarize our major contributions of this paper as follows:
@@
  • Empirically, we evaluated MARS-AdamW (or simply MARS) on GPT-2 fine-tuning tasks using the OpenWebText dataset. MARS demonstrated superior performance on GPT-2 large: it reached a validation loss of 2.58 within 27 billion tokens, whereas AdamW required 50 billion tokens to achieve the same level. The final validation loss achieved by MARS on GPT-2 large is 2.53, compared to 2.56 for AdamW. Furthermore, on the downstream task Hellaswag, MARS improved accuracy to 44.20%, outperforming AdamW’s 42.31% after training on 50 billion tokens, highlighting its superior performance in training large models.
@@ 3.1 MARS Framework
SuperAdam (Huang et al., 2021) also incorporates the STORM into the design of adaptive gradient methods. However, their precondition matrix can be viewed as a special case of our general framework. SuperAdam’s design focuses on diagonal precondition matrix and draws heavily from the design used in Adam (Kingma, 2014), AdaGrad-Norm (Ward et al., 2020), and AdaBelief (Zhuang et al., 2020). Furthermore, their preconditioner matrix is designed following Adam’s structure but does not account for the revised definition of variance-reduced momentum, resulting in a significant mismatch between the first-order and second-order momentum. We will further clarify these differences when discussing specific instances of our framework.
@@
$\mathbf{c}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}}\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}}_{\text{scaled gradient correction}}.$ (3.5)
When $\gamma_{t}=1$, (3.5) reduces to the second term of (3.3). On the other hand, when $\gamma_{t}=0$, (3.5) reduces to the stochastic gradient. Thus, $\mathbf{c}_{t}$ can be seen an gradient estimator with adjustable variance control. Following standard techniques in deep learning practice, we also perform gradient clipping on $\mathbf{c}_{t}$, which is calculated by:
@@
Even when using approximated Hessian matrices, second-order algorithms mentioned above remain more computationally intensive compared to first-order gradient updates. Thus, another line of research focuses on approximating the Hessian matrix through diagonal matrices, as seen in optimization algorithms like AdaGrad (Duchi et al., 2011), RMSProp (Tieleman, 2012), AdaDelta (Zeiler, 2012), Adam (Kingma, 2014) and AdamW (Loshchilov, 2017), etc. This approach to diagonal preconditioning effectively transforms the updates into a first-order method, assigning adaptive learning rates to each gradient coordinate. For example, in AdaGrad (Duchi et al., 2011), the preconditioned matrix is defined by:
@@ 3.2.1 MARS-AdamW
The first instance of MARS is built up on the idea of Adam/AdamW (Loshchilov, 2017). To automatically adjust the learning rate and accelerate convergence, Adam (Kingma, 2014) adopts the adaptive preconditioned gradient in (3.7) together with a bias correction and $\ell_{2}$ regularization. AdamW (Loshchilov, 2017) further changes the $\ell_{2}$ regularization to a decoupled weight decay. Overall, the full AdamW updates can be summarized as follows:
@@
In MARS-AdamW, we implement the preconditioned variance-reduced update as in (3.4), and utilize the same definitions for $\mathbf{H}_{t}$, $\epsilon$, and weight decay as those specified in AdamW. For $\mathbf{v}_{t}$, different from the EMA of squared gradients $\mathbf{g}_{t}^{2}$ in AdamW, we redefine it to fit our variance-reduced stochastic gradient. Specifically, we denote the summation of the stochastic gradient and the scaled gradient correction term by $\mathbf{c}_{t}$ and define $\mathbf{v}_{t}$ as the EMA of $\mathbf{c}_{t}^{2}$ as follows:
@@
Here, $\gamma_{t}$ is a scaling parameter parameter that controls the strength of gradient correction. When $\gamma_{t}=0$, the algorithm reduces to AdamW. Conversely, when $\gamma_{t}=1$, (3.14) aligns with the STORM momentum. Combining (3.13), (3.14), (3.15) together with (3.12) and the mirror descent update (3.2), we derive the MARS-AdamW algorithm in Algorithm 2. In practice, $\gamma_{t}$ is often set between 00 and $1$. Moreover, we employ gradient clipping-by-norm to $\mathbf{c}_{t}$ at Line 5, following the standard gradient clipping technique performed in neural network training.
@@
Connection with Adan.
@@
$\mathbf{y}_{t}$ $=\beta_{1}\mathbf{y}_{t-1}+(1-\beta_{1})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t}),$ (3.16)
$\mathbf{z}_{t}$ $=\beta_{2}\mathbf{z}_{t-1}+(1-\beta_{2})\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)},$ (3.17)
$\mathbf{m}_{t}$ $:=\mathbf{y}_{t}+\beta_{2}\mathbf{z}_{t}.$ (3.18)
@@
which is a special case of MARS-approx’s momentum with $\gamma_{t}=1-\beta_{1}$. It is worth noting that although motivated by the Nesterov’s momentum, Adan’s momentum updates in (3.16), (3.17) and (3.18) cannot recover Nesterov’s momentum unless $\beta_{1}=\beta_{2}$.
@@ 3.2.2 MARS-Lion
Connection with Lion.
Lemma 3.4.
For any sequence $\{\mathbf{g}_{t}\in\mathbb{R}^{d}\}_{t=0,1,\ldots}$, consider the following updates of $\mathbf{m}_{t}$ for any constant factors $a_{1},a_{2},b_{1}$, and $b_{2}$:
$\mathbf{u}_{t}$ $=a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{g}_{t},$ (3.23)
$\mathbf{m}_{t}$ $=b_{1}\mathbf{u}_{t}+b_{2}\mathbf{g}_{t}.$ (3.24)
The updates are equivalent to
$\mathbf{m}_{t}=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{g}_{t}+a_{1}b_{2}(\mathbf{g}_{t}-\mathbf{g}_{t-1}).$
Setting $\mathbf{g}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{1}$, $a_{2}=1-\beta_{1}$, $b_{1}=\beta_{2}$, $b_{2}=1-\beta_{2}$ in Lemma 3.4, and shifting the index of $\mathbf{u}_{t}$ by taking $\mathbf{u}_{t}=\mathbf{u}_{t-1}$, $\mathbf{u}_{t-1}=\mathbf{u}_{t-2}$ in Lemma 3.4, we can show that Lion momentum updates in (3.19) and (3.20) are equivalent to the following single momentum update:
@@ 3.2.3 MARS-Shampoo
Shampoo (Gupta et al., 2018) introduces a preconditioning approach that operates on the eigenspace of matrices. Given the gradient matrix $\mathbf{G}_{t}:=\nabla f_{t}(\mathbf{x}_{t},\bm{\xi}_{t})\in\mathbb{R}^{m\times n}$, the update rules of Shampoo are displayed as follows:
$\mathbf{L}_{t}$ $=\mathbf{L}_{t-1}+\mathbf{G}_{t}\mathbf{G}_{t}^{\top},$
$\mathbf{R}_{t}$ $=\mathbf{R}_{t-1}+\mathbf{G}_{t}^{\top}\mathbf{G}_{t},$
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\mathbf{L}_{t}^{-1/4}\mathbf{G}_{t}\mathbf{R}_{t}^{-1/4},$ (3.27)
where $\mathbf{x}_{t}\in\mathbb{R}^{m\times n}$ (slightly abusing notation) represents the corresponding weight matrix. It has been shown that the two-sided preconditioning in (3.27) is equivalent to preconditioning on the flattened vector $\mathbf{g}_{t}:=\text{vec}(\mathbf{G}_{t})$ with a Kronecker product (Gupta et al., 2018; Morwani et al., 2024)
$\mathbf{H}_{t}:=\Big{(}\sum_{\tau=1}^{t}\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\Big{)}^{1/4}\bigotimes\Big{(}\sum_{\tau=1}^{t}\mathbf{G}_{t}^{\top}\mathbf{G}_{t}\Big{)}^{1/4}.$
In practice, an exponential moving average (EMA) is often used in place of the direct summation. The update rule in (3.27) can be simplified to $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\big{(}\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\big{)}^{-1/4}\mathbf{G}_{t}\big{(}\mathbf{G}_{t}^{\top}\mathbf{G}_{t}\big{)}^{-1/4}$. This is equivalent to performing preconditioning on the eigenspace of $\mathbf{G}_{t}$:
$\mathbf{U}_{t},\mathbf{\Sigma}_{t},\mathbf{V}_{t}$ $=\text{SVD}(\mathbf{G}_{t}),$
@@
Connection with Muon. Muon (Jordan et al., 2024) is a recently proposed algorithm that utilizes the Newton-Schulz iteration (Higham, 2008; Schulz, 1933) to solve the SVD problem. It has demonstrated superior performance in terms of convergence speed when compared with AdamW and Shampoo in training large language models. The update rules of Muon are demonstrated as follows:
@@
Applying Lemma 3.4 to (3.29) and (3.30), with $\mathbf{g}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\mu$, $a_{2}=1$, $b_{1}=\mu$, $b_{2}=1$, we obtain an equivalent single update of momentum:

### v2

@@ Abstract
Training deep neural networks—and more recently, large models demands efficient and scalable optimizers. Adaptive gradient algorithms like Adam, AdamW, and their variants have been central to this task. Despite the development of numerous variance reduction algorithms in the past decade aimed at accelerating stochastic optimization in both convex and nonconvex settings, variance reduction has not found widespread success in training deep neural networks or large language models. Consequently, it has remained a less favored approach in modern AI. In this paper, to unleash the power of variance reduction for efficient training of large models, we propose a unified optimization framework, MARS (Make vAriance Reduction Shine), which reconciles preconditioned gradient methods with variance reduction via a scaled stochastic recursive momentum technique. Within our framework, we introduce three instances of MARS that leverage preconditioned gradient updates based on AdamW, Lion, and Shampoo, respectively. We also draw a connection between our algorithms and existing optimizers. Experimental results on training GPT-2 models indicate that MARS consistently outperforms AdamW by a large margin. The implementation of MARS is available at https://github.com/AGI-Arena/MARS.
@@ 1 Introduction
Adaptive gradient methods such as Adam (Kingma and Ba, 2015) and AdamW (Loshchilov and Hutter, 2019) have become the predominant optimization algorithms in deep learning. With the surge of large language models, the majority of the renowned models, including GPT-2 (Radford et al., 2019), GPT-3 (Brown, 2020), PaLM (Chowdhery et al., 2023) and Llama 3 (Dubey et al., 2024) are trained with adaptive gradient methods. Numerous efforts have been made to improve adaptive gradient methods from both first-order and second-order optimization perspectives. For example, You et al. (2019) introduced LAMB, a layerwise adaptation technique that boosts training efficiency for BERT (Devlin, 2018). Using symbolic search, Chen et al. (2023) developed Lion, achieving faster training and reduced memory usage. Liu et al. (2023) designed Sophia, leveraging stochastic diagonal Hessian estimators to accelerate training. Gupta et al. (2018) proposed Shampoo, which performs stochastic optimization over tensor spaces with preconditioning matrices for each dimension. Anil et al. (2020) further refined Shampoo to give a scalable, practical version. Recently, Vyas et al. (2024) showed that Shampoo is equivalent to Adafactor (Shazeer and Stern, 2018) in the eigenbasis of Shampoo’s preconditioner, and introduced SOAP, which stabilizes Shampoo with Adam. Importantly, recent studies (Kaddour et al., 2024; Zhao et al., 2024) have shown that these optimizers perform on par with AdamW in LLM pretraining, yet do not outperform it. This suggests the ongoing challenge in developing adaptive gradient methods superior to Adam and AdamW for large-scale model training.
Since adaptive gradient methods face challenges of high stochastic gradient variance, and language model training inherently involves a high-variance optimization problem (McCandlish et al., 2018), it is natural to consider variance reduction techniques to address this challenge. There exists a large body of literature on variance reduction for stochastic optimization, such as SAG (Roux et al., 2012), SVRG (Johnson and Zhang, 2013), STORM (Cutkosky and Orabona, 2019), which can improve the convergence of stochastic optimization. However, variance reduction has not found widespread success in training deep neural networks or large language models. Defazio and Bottou (2019) discussed why variance reduction can be ineffective in deep learning due to factors such as data augmentation, batch normalization, and dropout, which disrupt the finite-sum structure required by variance reduction principles. Nevertheless, in training language models, data augmentation, batch normalization, and dropout are nowadays rarely used, which opens the door for applying variance reduction techniques in optimizing these models. This naturally leads to the following research question:
@@
In this paper, we answer the above question affirmatively by introducing a novel optimization framework called MARS (Make vAriance Reduction Shine), which incorporates variance reduction into adaptive gradient methods. Notably, we introduce a scaling parameter $\gamma\in(0,1]$ into the stochastic recursive momentum (STORM) (Cutkosky and Orabona, 2019) to adjust the strength of variance reduction and define a new gradient estimator. This gradient estimator undergoes gradient clipping and is subsequently subjected to exponential averaging. When the scaling parameter is set to $\gamma=1$, it recovers the vanilla STORM momentum. In addition, the second-order momentum update is defined by the reweighted intermediate variable. These together ensure optimization stability throughout the training process. We summarize our major contributions of this paper as follows:
@@
  • Theoretically, we prove that MARS achieves an $\mathcal{O}(T^{-2/3})$ convergence rate to the first-order stationary points, which improves upon the $\mathcal{O}(T^{-1/2})$ convergence rate of Adam/AdamW  (Zhou et al., 2018; Chen et al., 2018b; Huang et al., 2021). Here $T$ is the number of iterations. Moreover, we precisely characterize the dependence of the upper bound on the scaling parameter $\gamma$, which controls the strength of variance reduction. When $\gamma=1$, the proposed momentum reduces to the STORM momentum, and the convergence rate recovers that of SuperAdam (Huang et al., 2021). Notably, an optimal $\gamma$ schedule exists that enables a faster convergence rate, demonstrating that the flexibility in choosing $\gamma$ leads to a tighter upper bound.
  • Empirically, we evaluated MARS on GPT-2 fine-tuning tasks using the OpenWebText dataset. It demonstrates superior performance on GPT-2 large: AdamW requires 50 billion tokens to reaches a validation loss of 2.58, whereas MARS only requires 28 billion tokens, and it achieves a final validation loss of 2.51. Furthermore, on the downstream task Hellaswag, MARS improved accuracy to 44.64%, outperforming AdamW’s 41.70% after training on 50 billion tokens.
@@ 3.1 MARS Framework
SuperAdam (Huang et al., 2021) also incorporates the STORM into the design of adaptive gradient methods. However, their precondition matrix can be viewed as a special case of our general framework. SuperAdam’s design focuses on diagonal precondition matrix and draws heavily from the design used in Adam (Kingma and Ba, 2015), AdaGrad-Norm (Ward et al., 2020), and AdaBelief (Zhuang et al., 2020). Furthermore, their preconditioner matrix is designed following Adam’s structure but does not account for the revised definition of variance-reduced momentum, resulting in a significant mismatch between the first-order and second-order momentum. We will further clarify these differences when discussing specific instances of our framework.
@@
$\mathbf{c}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}}\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}}_{\text{scaled gradient correction}}.$
When $\gamma_{t}=1$, the above reduces to the second term of (3.3). On the other hand, when $\gamma_{t}=0$, it reduces to the stochastic gradient. Thus, $\mathbf{c}_{t}$ can be seen a gradient estimator with adjustable variance control.
Following standard techniques in deep learning practice, we also perform gradient clipping on $\mathbf{c}_{t}$, which is calculated by:
@@
Why $\gamma_{t}$ improves convergence
A similar idea of adjusting variance reduction strength has been proposed in Yin et al. (2023) for SVRG. For the STORM updates used in our work, we provide an intuition by the estimation of
$\mathbb{E}[X+(\gamma Y-\mathbb{E}Y)]^{2},$
where $X$ and $Y$ are two independent random variables. [^1] In our case, $Y=\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})$. See Proof of Lemma B.2 for details.
  • When $\gamma=1$, this simplifies to $\mathbb{E}[X+(Y-\mathbb{E}Y)]^{2}=\mathbb{E}X^{2}+\mathrm{Var}(Y)$.
  • When $\gamma=0$ it reduces to $\mathbb{E}[X-\mathbb{E}Y]^{2}$, which can be significantly larger than choosing $\gamma=1$ when $\mathbb{E}XY<0$.
  • There exists an optimal choice of $\gamma^{*}$, given by $\gamma^{*}=1-\frac{\mathbb{E}XY+\mathrm{Var}(Y)}{\mathbb{E}Y^{2}}$, such that
    $\mathbb{E}[X+(\gamma Y-\mathbb{E}Y)]^{2}$ $=\mathbb{E}[X+(Y-\mathbb{E}Y)]^{2}-\big{(}\mathbb{E}XY+\mathrm{Var}(Y)\big{)}^{2}/\mathbb{E}Y^{2},$
which yields a smaller variance compared to any other choice of $\gamma$. We provide a convergence analysis of Algorithm 1 in Theorem 4.4 in Section 4. We prove that under standard assumptions, MARS achieves a superior convergence rate of $\mathcal{O}(T^{-2/3})$, outperforming the $\mathcal{O}(T^{-1/2})$ rate attainable by AdamW and its variants (Zhou et al., 2018; Chen et al., 2018b; Huang et al., 2021).
@@
Even when using approximated Hessian matrices, second-order algorithms mentioned above remain more computationally intensive compared to first-order gradient updates. Thus, another line of research focuses on approximating the Hessian matrix through diagonal matrices, as seen in optimization algorithms like AdaGrad (Duchi et al., 2011), RMSProp (Tieleman, 2012), AdaDelta (Zeiler, 2012), Adam (Kingma and Ba, 2015) and AdamW (Loshchilov and Hutter, 2019), etc. This approach to diagonal preconditioning effectively transforms the updates into a first-order method, assigning adaptive learning rates to each gradient coordinate. For example, in AdaGrad (Duchi et al., 2011), the preconditioned matrix is defined by:
@@ 3.2.1 MARS-AdamW
The first instance of MARS is built up on the idea of Adam/AdamW (Loshchilov and Hutter, 2019). To automatically adjust the learning rate and accelerate convergence, Adam (Kingma and Ba, 2015) adopts the adaptive preconditioned gradient in (3.6) together with a bias correction and $\ell_{2}$ regularization. AdamW (Loshchilov and Hutter, 2019) further changes the $\ell_{2}$ regularization to a decoupled weight decay. Overall, the full AdamW updates can be summarized as follows:
@@
In MARS-AdamW, we implement the preconditioned variance-reduced update as in (3.4), and utilize the same definitions for $\mathbf{H}_{t}$, $\epsilon$, and weight decay as those specified in AdamW. For $\mathbf{v}_{t}$, different from the EMA of squared gradients $\mathbf{m}_{t}^{2}$ in AdamW, we redefine it to fit our variance-reduced stochastic gradient. Specifically, we denote the summation of the stochastic gradient and the scaled gradient correction term by $\mathbf{c}_{t}$ and define $\mathbf{v}_{t}$ as the EMA of $\mathbf{c}_{t}^{2}$ as follows:
@@
Here, $\gamma_{t}$ is a scaling parameter that controls the strength of gradient correction. When $\gamma_{t}=0$, the algorithm reduces to AdamW. Conversely, when $\gamma_{t}=1$, (3.13) aligns with the STORM momentum. Combining (3.12), (3.13), (3.14) together with (3.11) and the mirror descent update (3.2), we derive the MARS-AdamW algorithm in Algorithm 2. In practice, $\gamma_{t}$ is often set between 00 and $1$. Moreover, we employ gradient clipping-by-norm to $\mathbf{c}_{t}$ at Line 5, following the standard gradient clipping technique performed in neural network training. We provide a convergence analysis of Algorithm 2 in Theorem 4.6 in Appendix 4.
@@
Connection between MARS-AdamW and Adan.
@@
$\mathbf{y}_{t}$ $=\beta_{1}\mathbf{y}_{t-1}+(1-\beta_{1})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t}),$
$\mathbf{z}_{t}$ $=\beta_{2}\mathbf{z}_{t-1}+(1-\beta_{2})\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)},$
$\mathbf{m}_{t}$ $:=\mathbf{y}_{t}+\beta_{2}\mathbf{z}_{t}.$
@@
which is a special case of MARS-approx’s momentum with $\gamma_{t}=1-\beta_{1}$. It is worth noting that although motivated by the Nesterov’s momentum, Adan’s momentum updates cannot recover Nesterov’s momentum unless $\beta_{1}=\beta_{2}$.
@@ 3.2.2 MARS-Lion
Setting $\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{1}$, $a_{2}=1-\beta_{1}$, $b_{1}=\beta_{2}$, $b_{2}=1-\beta_{2}$ in Lemma C.1, and shifting the index of $\mathbf{u}_{t}$ by taking $\mathbf{u}_{t}=\mathbf{u}_{t-1}$, $\mathbf{u}_{t-1}=\mathbf{u}_{t-2}$ in Lemma C.1, we can show that Lion momentum updates in (3.15) and (3.16) are equivalent to the following single momentum update:
@@ 3.2.3 MARS-Shampoo
Shampoo (Gupta et al., 2018) introduces a preconditioning approach that operates on the eigenspace of matrices. Given the gradient matrix $\mathbf{m}_{t}:=\nabla f_{t}(\mathbf{x}_{t},\bm{\xi}_{t})\in\mathbb{R}^{m\times n}$, the update rules of Shampoo are displayed as follows:
$\mathbf{L}_{t}$ $=\mathbf{L}_{t-1}+\mathbf{m}_{t}\mathbf{m}_{t}^{\top},$
$\mathbf{R}_{t}$ $=\mathbf{R}_{t-1}+\mathbf{m}_{t}^{\top}\mathbf{m}_{t},$
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\mathbf{L}_{T}^{-1/2}\mathbf{m}_{t}\mathbf{R}_{T}^{-1/2},$ (3.21)
where $\mathbf{x}_{t}\in\mathbb{R}^{m\times n}$ (slightly abusing notation) represents the corresponding weight matrix. It has been shown that the two-sided preconditioning in (3.21) is equivalent to preconditioning on the flattened vector $\mathbf{m}_{t}:=\text{vec}(\mathbf{m}_{t})$ with a Kronecker product (Gupta et al., 2018; Morwani et al., 2024)
$\mathbf{H}_{t}:=\Big{(}\sum_{\tau=1}^{t}\mathbf{G}_{\tau}\mathbf{G}_{\tau}^{\top}\Big{)}^{1/4}\otimes\Big{(}\sum_{\tau=1}^{t}\mathbf{G}_{\tau}^{\top}\mathbf{G}_{\tau}\Big{)}^{1/4}.$
In practice, an exponential moving average (EMA) is often used in place of the direct summation. The update rule in (3.21) can be simplified to $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\big{(}\mathbf{m}_{t}\mathbf{m}_{t}^{\top}\big{)}^{-1/4}\mathbf{m}_{t}\big{(}\mathbf{m}_{t}^{\top}\mathbf{m}_{t}\big{)}^{-1/4}$. This is equivalent to performing preconditioning on the eigenspace of $\mathbf{m}_{t}$:
$\mathbf{U}_{t},\mathbf{\Sigma}_{t},\mathbf{V}_{t}$ $=\text{SVD}(\mathbf{m}_{t}),$
@@
Connection between MARS-Shampoo and Muon.
Muon (Jordan et al., 2024) is a recently proposed algorithm that utilizes the Newton-Schulz iteration (Higham, 2008; Schulz, 1933) to solve the SVD problem. It has demonstrated superior performance in terms of convergence speed when compared with AdamW and Shampoo in training large language models. The update rules of Muon are demonstrated as follows:
@@
Applying Lemma C.1 to (3.23) and (3.24), with $\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\mu$, $a_{2}=1$, $b_{1}=\mu$, $b_{2}=1$, we obtain an equivalent single update of momentum: